OpenStudy (anonymous):

How to find Inverse of f(x) = 2x + 3/4 ?

OpenStudy (amistre64):

switch your x and y then sove for "x" like this: y = 2x + (3/4) switch em x = 2y + (3/4) ... now solve for "y" x - (3/4) = 2y x/2 - 3/8 = y

OpenStudy (anonymous):

I always like the nest approach: \[f(f^{-1}(x)) = x \text{, by the definition of the inverse}\] \[\implies 2f^{-1}(x) + 3/4 = x\] \[\implies 2f^{-1}(x) = x-3/4\] \[\implies f^{-1}(x) = x/2-3/8\]